OpenStudy (anonymous):

Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (anonymous):

the one that is called the "discriminant" will tell you.... see if you can google that...

OpenStudy (anonymous):



OpenStudy (anonymous):

das da one ^^^^

OpenStudy (anonymous):

Use the given equation and plug it into the discriminant and your answer will tell you if you can factor or not.

OpenStudy (anonymous):

I dont understand how to get the answer though.

OpenStudy (anonymous):

The discriminant is b^2-4ac. Therefore, using the given equation..2x^2 +7x +3=0 You plug in the a, b, c. a=2 b=7 c=3 (7)^2 - 4(2)(3)= 49-48=1

OpenStudy (anonymous):

so whats the answer you have to type into the box?